(1) If a,b,c are real numbers uch that a^2+2b=6, b^2+4c=-7, c^2+6a=-13, then the value of a^2+b^2+c^2 is equal to:
(A) 14 (B) 21 (C) 28 (D) 35

Answer by Edwin McCravy(20086)

If a,b,c are real numbers uch that a^2+2b=6, b^2+4c=-7, c^2+6a=-13, then the value of a^2+b^2+c^2 is equal to:
(A) 14 (B) 21 (C) 28 (D) 35
There is no solution.